OpenStudy (anonymous):

i don't really get what to do. "Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. Find the two numbers."

OpenStudy (anonymous):

x+y=170 x-y=26

OpenStudy (anonymous):

x + y =170 x - y = 26 2x = 196 x = 98 y = 72

OpenStudy (anonymous):

x+y=98 and 72

OpenStudy (anonymous):

okay, thank you :)